### v3

# WonderVerse: Extendable 3D Scene Generation with Video Generative Models

Abstract
We introduce WonderVerse, a simple but effective framework for generating extendable 3D scenes. Unlike existing methods that rely on iterative depth estimation and image inpainting, often leading to geometric distortions and inconsistencies, WonderVerse leverages the powerful world-level priors embedded within video generative foundation models to create highly immersive and geometrically coherent 3D environments. Furthermore, we propose a new technique for controllable 3D scene extension to substantially increase the scale of the generated environments. Besides, we introduce a novel abnormal sequence detection module that utilizes camera trajectory to address geometric inconsistency in the generated videos. Finally, WonderVerse is compatible with various 3D reconstruction methods, allowing both efficient and high-quality generation. Extensive experiments on 3D scene generation demonstrate that our WonderVerse, with an elegant and simple pipeline, delivers extendable and highly-realistic 3D scenes, markedly outperforming existing works that rely on more complex architectures.
Figure 1: WonderVerse is able to create large-scale, coherent, extendable, and high-quality 3D scenes from a text.
1 Introduction
3D scene generation is a fundamental task in the vision community due to its wide-ranging applications in Virtual Reality, Mixed Reality (VR/MR), robotics, self-driving vehicles, and more. However, creating extendable and large-scale 3D scenes poses significant challenges. First, there is a lack of large-scale, high-quality datasets of 3D scenes. Second, generating realistic and geometrically accurate 3D scenes is inherently complex due to the unstructured nature of 3D space and the need to capture intricate object relationships and environmental context. Third, seamlessly extending 3D scenes while maintaining consistency and coherence remains a significant challenge.
To generate extendable 3D scenes, recent studies [20, 14, 9, 56, 55] leverage 2D generative models for 3D scene generation. These methods typically iteratively extends a given image to synthesize views from new camera poses. Typically, they first estimate a depth map from the input image, project it into 3D, and then use an image inpainting model to extend the scene from novel viewpoints. This iterative process builds and extends the 3D environment. However, the quality of scenes generated by these approaches remains limited. As shown in Figure 4, a primary issue is geometric distortion, arising from the scale ambiguity of single-view depth estimation and the lack of consistent camera parameters across iterations. Additionally, error accumulation is a concern due to the iterative application of depth estimation and inpainting during scene expansion. Furthermore, seams and discontinuities can appear between regions generated in different iterations. Overall, the iterative discrete-view-based generation pipeline inherently results in unsatisfactory scene quality.
Going beyond existing works that rely on image generation technics, we propose a simple yet effective approach to leverage the recent advancements of video generative models [23, 7, 42, 58, 61, 53] to produce highly immersive and superior-quality 3D scenes. Our approach is motivated by an intuitive but thought-provoking fact that a video with a circular camera trajectory can naturally represent a 3D scene with good view consistency. This implies that the video, compared with images, has a smaller domain gap with 3D representations. In addition, we can fully utilize the rich prior knowledge of video generative models for 3D scene generation without requiring any 3D scene dataset. More importantly, the generation process does not require iteratively generating discrete view of images in previous methods; thus, unlike existing works, our approach is inherently free from issues like scale ambiguity of depth, camera parameter inconsistency, and error accumulation throughout iterative generation.
To achieve this, a natural initial thought is to combine video generative models with 3DGS (3D Gaussian Splatting) [22] for 3D scene generation. Nevertheless, this simple combination does not yield satisfactory results and cannot enable extendable 3D scene generation. Firstly, the scale of the generated scene is restricted by the capabilities of the video generative model itself. Moreover, as recent works [33, 6] have shown, generated videos can suffer from geometric inconsistencies across different frames. These inconsistencies result in substantial distortions in the resulting 3D scenes.
To address these issues, we present WonderVerse, a simple but effective extendable 3D scene generative model powered by video generative models. After generating a circular video sequence from the text input, we present a new 3D scene extension approach by extending the video at different views to form multiple extended videos, enabling large-scale 3D scene generation. Besides, we find COLMAP camera pose estimation effectively flags the geometric issue: inconsistent videos produce erratic camera poses and discontinuous trajectory. Based on the observation, we introduce an abnormal sequence detection module that evaluates the continuity of the estimated camera trajectory to enhance geometric coherence. With the consistent and coherent video sequences, we reconstruct the 3D scene. Note that our approach is compatible to various 3D reconstruction approaches for both efficient (DUSt3R [43]) and high-quality (3DGS [22]) 3D scene generation.
As illustrated in Figure 1, our WonderVerse can create a extendable and highly immersive 3D scene from a piece of text. Furthermore, both qualitative and quantitative experimental results demonstrates that WonderVerse’s neat and simple pipeline leads to state-of-the-art extendable and highly-realistic extendable 3D scene generation. Our approach demonstrates that sophisticated results can be attained through our highly elegant design. Codes will be released upon publication.
In summary, the key contributions of WonderVerse are:
  • We propose a simple yet effective approach to leverage video generative models for the extendable 3D scene generation task, surpassing image-based iterative pipelines.
  • We introduce a new video-driven 3D scene extension approach to scale up the generated environments.
  • We develop an abnormal sequence detection module to enhance the geometric consistency of both generated videos and 3D scenes.
  • WonderVerse achieves state-of-the-art extendable and highly-realistic 3D scene generation through a surprisingly neat and simple pipeline, significantly outperforming existing methods with far more complex architectures.
2 Related Works
In this section, we will briefly introduce the development of novel view synthesis (NVS), video generation for 3D reconstruction, and extendable 3D scene generation.
2.1 Novel View Synthesis
Novel view synthesis (NVS) is a fundamental vision task to produce a 3D scene from multi-view input images, allowing the rendering of images from novel views. Some traditional methods tackle the problem by formulating 3D scenes as light fields [10, 24], multi-view depth images [16], or blending of a collection of input images [40]. Rather than relying on handcrafted heuristics for constructing 3D scenes, the community later adopts deep learning techniques for NVS [18, 31]. Among these approaches, NeRF [31] has achieved success in improving the image’s quality, resulting in an explosion of follow-up methods [2, barron2021mip, 59, 4, 35, 30] Afterwords, 3D Gaussian Splatting (3DGS) [22] is introduced as a 3D representation that enables fast training and real-time rendering while maintaining high-quality reconstruction results. This has motivated a series of follow-up works [21, 57, 11] that aim to further improve the quality and efficiency of the 3DGS representation. Recently, instead of employing an expensive optimization process, DUSt3R [43] innovatively introduces a pointmap representation, enabling end-to-end 3D reconstruction with fast speed and promising performance. In this work, we demonstrate that our approach is compatible to different 3D reconstruction methods, such as 3DGS and DUSt3R.
2.2 Image and Video Generative Models for MVS and 3D Reconstruction
Recent progress in foundation models for image [34] and video [blattmann2023stable] generation has driven the exploration of their use in 3D reconstruction. Image generative models can be fine-tuned with multi-view data for novel-view synthesis and 3D reconstruction of objects [27, 29] and scenes [15, 48]. This strategy has been extended to video foundation models, enabling video generation with controlled camera motion for 3D reconstruction [1, 42, 25, 41, 52, 17, 44]. Nevertheless, these methods primarily target object-level or limited-scale reconstruction and cannot directly adapt to our extendable 3D scene generation task. Our objective is to generate extendable 3D scenes, requiring models with the imaginative ability to extend 3D scenes beyond the initial input.
2.3 Extendable 3D Scene Generation
Another branch of works explore extendable 3D scene generation. Existing works like [50, 51, 60, 12, 26] only focus on urban environments and cannot be generalized to other scenes, restricting their semantic scope. For unbounded nature scenes, methods like [8, 5] requires multi-view data, while others [49, 28, 45] rely on 3D scene datasets. To avoid such data dependencies, Text2Room [20], SceneScape [14], LucidDreamer [9], WonderJourney [56], WonderWorld [55] generate scenes iteratively with depth estimation and inpainting. Among existing approaches, these methods are the most relevant ones to our work. However, their iterative pipelines inherently suffers from geometric distortion, discontinuity, and error accumulation during scene expansion, ultimately hampering scene quality and motivating our WonderVerse framework. In this work, we seek to overcome the above challenges and leverage video generative models to generate more immersive and expandable scenes.
3 Methods
In this section, we will introduce our WonderVerse which is illustrated in Figure 2. Given a text description, the WonderVerse first generates a video that continuously presents the scene circularly. It then extends the scene by creating videos guided by the left and right views. Additionally, the framework estimates the camera pose sequence based on the generated videos and detects abnormal sequences by identifying discontinuities in camera movements. The frames corresponding to these abnormal sequences will be regenerated until all pass the filtering criteria. Finally, a 3D scene reconstruction is performed to construct the complete scene.
Figure 2: Illustration of our WonderVerse. This framework includes: (a) a text-guided video generation and extension module that produces a video of a scene circularly in a continuous shot, followed by extensions to both sides; (b) a camera parameter estimation module that predicts the camera pose sequence; (c) a abnormal sequence detection module that identifies discontinuous camera poses and regenerates the corresponding videos; and (d) a 3D scene reconstruction and rendering module to construct the generated scene.
3.1 Video Generation and Extension
Most existing works on extendable 3D scene generation relies on iterative image generation. Such a design can hinder the quality of the generated results due to the significant domain gap between images and 3D representations. Simply stitching multiple images together to construct a scene often leads to geometric inconsistencies. Inspired by the rapid development of video generative models like [23], we reconsider the problem by leveraging superior generative capability of such models. Therefore, we propose to address the extendable 3D scene generation task through the lens of video generation.
Given Text Prompt $P$, we perform text-guided video generation [23] to obtain high-quality initial video $V_{init}\in\mathbb{R}^{T\times H\times W\times 3}$, where $H$ and $W$ denotes the hight and weight of each frames, respectively. $T$ denotes the length of the video. $3$ denotes RGB channels. To enrich the scene, we further perform video-based scene extensions by extending the initial video. To maintain style consistency, we use the first frame $V_{init}^{t=1}$ and last frame $V_{init}^{t=T}$ of the initial video as references and generate two new videos from left and right, respectively: $V_{extend}^{left}\in\mathbb{R}^{T^{\prime}\times H\times W\times 3}$ and $V_{extend}^{right}\in\mathbb{R}^{T^{\prime}\times H\times W\times 3}$ with prompts different from the initial input, $V_{init}$. Then, we combine them into an extended video $\{V_{extend}^{left},V_{init},V_{extend}^{right}\}$ with length $T+2T^{\prime}$. The video extension can be applied repeatedly $n$ times, taking $V_{input}$ as the initial video for the next iteration, further enriching the scene and yielding $V_{input}=\{V_{extend}^{left_{n}},V_{extend}^{left_{n-1}},...,V_{init},...,V_{extend}^{right_{n-1}},V_{extend}^{right_{n}}\}=\{V_{1},V_{2},...,V_{2n+1}\}$. Figure 2 (a) illustrates two iterations of video extension. More details can be found in Sec. 4.
3.2 Camera Parameters Estimation
Previous works [20, 9, 56, 55] extend images to 3D scenes through iterative image generation, camera parameter estimation, and rendering. However, their camera trajectories are often subjectively defined, leading to potential angle mismatches between rendered and synthesized images. This results in diminished scene quality and visual inconsistencies. In contrast, leveraging the inherent continuity of video, our approach uses COLMAP [38, 39] to compute consistent camera parameters for 3D scene optimization, thus achieving enhanced geometric consistency in the generated 3D scenes.
To acquire point clouds and camera poses, we utilize COLMAP, a Structure from Motion (SfM) and Multi-View Stereo (MVS) pipeline that yields accurate results for 3D scene initialization. For video data, we employ sequential matching as our feature-matching strategy within SfM to reduce computational complexity and accelerate estimation. Furthermore, to enhance matching quality, we apply guided matching, which leverages known geometric information to direct feature matching and constrain the search space.
3.3 Abnormal Sequence Detection
1Inputs: video length $l_{\text{video}}$,
2 input videos $V_{\text{input}}=\{V_{1},V_{2},...,V_{j},...,V_{2n+1}\}$,
3 video re-generation function $\Psi(\cdot)$,
4 camera extrinsic parameters list
5 $L=\{(R_{1},P_{1}),(R_{2},P_{2}),...,(R_{l},P_{l}),...\}$,
6 camera parameter estimation function $\Gamma(\cdot)$,
7 extrinsic parameter threshold $\theta=\{\theta_{R},\theta_{P}\}$
8Outputs: stable scene video $V_{\text{output}}=\{\tilde{V}_{1},\tilde{V}_{2},...,\tilde{V}_{j},...,\tilde{V}_{2n+1}\}$
9Initialization: sort $L$ by timestamp
10 $l\leftarrow 2$; $V_{\text{output}}\leftarrow V_{\text{input}}$
11while $l<|L|$ do
12       $\bigtriangleup{R}_{l},\bigtriangleup{P}_{l}\leftarrow\text{Abs}(R_{l}-R_{l+1}),\text{Abs}(P_{l}-P_{l+1})$
13      if $\bigtriangleup{R}_{l}>\theta_{R}$ or $\bigtriangleup{P}_{l}>\theta_{P}$ then
14             $j\leftarrow\text{GetCurrentVideoID}(l)$
15             $\tilde{V}_{j}\leftarrow\Psi(V_{j})$triangleright re-generate unstable segment
16             $V_{\text{output}}\leftarrow\text{Update}(V_{\text{output}},\tilde{V}_{j})$ $L\leftarrow\text{Update}(L,\Gamma(\tilde{V}_{j}))$
17             $\triangleright$ Re-estimate the camera parameters $l\leftarrow\text{GetStartFrameID}(\tilde{V}_{j})$
18             $\triangleright$ Back to the start of $\tilde{V}_{j}$
19      else
20             $l\leftarrow l+1$
21       end if
22
23 end while
return $V_{\text{output}}$
Algorithm 1 Abnormal sequence detection.
Although the video generation model is powerful, not all videos generated from prompts are suitable for scene generation. A common issue degrading 3D scene generation is the geometric inconsistency of video generative models [6, 33]. We observed that geometric inconsistencies are effectively revealed by erratic camera pose estimations from COLMAP. Assuming smooth video generation implies smooth camera motion, geometric inconsistencies disrupt pose estimation, resulting in random, inaccurate poses. We empirically found a strong correlation between visually identified geometric inconsistency and unreliable COLMAP poses across a dataset with 320 videos, indicating pose estimation as a practical indicator.
Based on the above observation, we design an abnormal sequence detection scheme that identifies these videos by locating discontinuities in the camera pose trajectory. We identify discontinuities by monitoring changes in camera position and rotation. A sequence is flagged as discontinuous and geometrically inconsistent if position shifts exceed 5 units or rotation changes exceed 0.5 degrees. If any abnormal sequences are detected, the corresponding videos undergo a re-generation scheme in which the video extension or generation, and the camera parameter estimation are performed again until all sequences pass the criteria.
Using the input videos and extracted camera poses (denoted by rotation R and position P for each frame), our algorithm iteratively evaluates the continuity of camera movements across video frames. In each iteration, the difference in camera poses between consecutive frames is computed. Should this difference surpass a defined threshold, the corresponding segment is marked for replacement, and camera parameters are re-estimated to ensure consistency. Conversely, if the difference remains within the threshold, the segment is deemed satisfactory, and the algorithm proceeds to the next frame. This iterative evaluation across all segments, allowing for the identification and correction of any anomalies in the video sequence.
3.4 3D Scene Reconstruction and Rendering
To obtain an immersive and high-quality scene, we employ an advanced 3D representations method, i.e. 3DGS [22], which supports real-time rendering and is faster in training.
Let us define a 3D Gaussian $i$ that is parameterized by a centroid $\mu_{i}$, a covariance matrix $\Sigma_{i}$, opacity $\sigma_{i}$ and color $c_{i}$, representing the first three degrees of the spherical harmonic (SH) coefficients. These properties are learnable and optimized during training. In practice, the covariance matrix is decomposed into a scaling matrix $S_{i}$ and a rotation matrix $R_{i}$ to ensure that the covariance matrix remains positive semi-definite and retains physical meaning, i.e. $\Sigma_{i}=R_{i}S_{i}S_{i}^{T}R_{i}^{T}$. To project 3D onto 2D, the covariance matrix is approximated by the Jacobian matrix and the world-to-camera matrix: $\Sigma’_{i}=JW\Sigma_{i}W^{T}J^{T}$. Then, the opacity of Gaussian at image plane $p^{\prime}_{i}$ is computed as follows:
$\alpha^{\prime}_{i}=\sigma_{i}\cdot exp(-\frac{1}{2}(p^{\prime}_{i}-\mu^{\prime}_{i})^{T}(\Sigma^{\prime}_{i})^{-1}(p^{\prime}_{i}-\mu^{\prime}_{i})).$ (1)
Finally, the color $C$ of pixel $p$ is:
$C(p)=\sum\limits_{i\in N}c_{i}\cdot\alpha^{\prime}_{i}\prod\limits_{j=1}^{i-1}(1-\alpha^{\prime}_{i}).$ (2)
After abnormal sequence detection, we derive a set of valid camera parameters, i.e. $L=\{(R_{1},P_{1}),(R_{2},P_{2}),...\}$, where $R,P$ denotes rotation and position respectively. Then we render $I_{ren}$ with 3DGS from each camera view, and it corresponds to a frame $I_{gt}$ of the generated video. Then we compute $\mathcal{L}_{1}$ loss and $\mathcal{L}_{D-SSIM}$ loss between $I_{ren}$ and $I_{gt}$. Finally, our loss function is:
$\mathcal{L}=(1-\lambda)\mathcal{L}_{1}+\lambda\mathcal{L}_{D-SSIM}.$ (3)
where $\mathcal{L}_{D-SSIM}$ is the SSIM loss that optimizes the image quality, and $\mathcal{L}_{1}$ is the L1 loss, $\lambda$ is set as 0.2.
In contrast to prior works that rely on point clouds [56, 20, 9] or less common data formats such as layered images [55] as intermediate representations, WonderVerse directly generates images and corresponding camera poses. This design significantly enhances compatibility, allowing us to leverage well-established 3D reconstruction methods that natively accept image and pose inputs, such as the recent DUSt3R [43]. This streamlined approach enables efficient 3D scene generation, with the potential for further improvements as reconstruction techniques advance.
4 Experiments
Figure 3: WonderVerse generates large-scale, extendable, cohenrent, and high-fidelity 3D scenes, both indoors and outdoors. Dashed lines show the camera’s direction during scene extension.
In this section, first, we will introduce the implementation details and evaluation metrics. Then we show our generative results and also compare with existing works both quantitatively and qualitatively. Subsequently, we conduct ablation study on our key modules. Note that we provide additional generative results in the supplementary materiel.
4.1 Experimental Settings
Implementation Details Starting with a text prompt, WonderVerse first generates an initial circular video. The first and last frames of this video serve as seeds for scene extension. To extend the scene, these seed frames are incorporated into the video generative model with text prompts, guiding the generation of additional videos that seamlessly extend the existing scene. This extension process is iterative and can be repeated for virtually infinite scene growth. In our experiments, we performed four extension iterations, adding two videos to the left and two to the right of the original scene. Our text prompts follow a template: “aerial shot, soft lighting, around left [or right], realistic, high-quality, displaying [scene description]”. Camera motion control was facilitated via the API provided by the video generative models. For initial scene generation, we utilized the Hunyuan video model [23]. However, as image-conditioned video generation was not yet available for Hunyuan at the time, we adopted Gen-3 Alpha [36] for the subsequent scene extension steps.
In our experiments, we generate an initial scene for each test example and extend it by creating four additional scenes, yielding five scenes per example and twenty in total. In contrast to WonderWorld’s image-based scenes, each extension in our approach produces a complete video scene. This video representation offers a significant advantage in immersiveness and richness over static image scenes of existing works.
Evaluation metrics Following recent work [55], we evaluate our scenes using CLIP-Score (CLIP-S) [19], Q-Align [47], and NIQE [32]. CLIP-S measures text-image alignment by comparing CLIP feature embeddings. Q-Align, trained with human scoring patterns and instruction tuning, assesses quality, correlating well with subjective human judgments. NIQE, a no-reference image quality metric, evaluates quality degradation by comparing spatial features to a pristine natural image model.
Figure 4: Qualitative comparison with existing works.
4.2 Our Results of Extendable 3D Scenes
In Figure 3, we present compelling examples of extendable indoor and outdoor 3D scenes generated by WonderVerse across. Figure 3 (a, b) illustrate the remarkable realism and rich detail achieved in our indoor scenes, with flawless geometric integrity. Also, the scenes are faithful and immersive visualizations of the described environments. For example, in (b) we generate a workshop, with “tools hanging on the wall”, a “sturdy workbench”, and “shelves stocked with materials”. All the objects are rendered with photorealistic quality and placed within a geometrically coherent space. Figure 3 (c, d) further exemplifies this capability in outdoor scenarios. For example, in (c), a playground in a residential area is generated, where all textually described components like “car” and “tree” are seamlessly integrated with remarkable visual fidelity. Also, our scenes have great geometric consistency and no obvious mistakes or breaks that look unnatural. This shows the strength of WonderVerse in making highly immersive and extendable 3D scenes. Please refer to the supplementary material for additional generative results across diverse scenes and styles of WonderVerse. Besides, we also demonstrate users can generate extendable 3D scenes iteratively using WonderVerse.
4.3 Comparison with Existing Works
Baseline Approaches We compare our WonderVerse with recent works WonderWorld [55] and LucidDreamer [9] both qualitatively and quantitatively. Since these two baselines need an image as input while our method only needs a text prompt as input, for fair comparison, we take a random frame of our generated video and feed it into their models for 3D scene generation. Note that WonderWorld [55]’s layered image representation, including a sky layer, makes it less suited for indoor scene generation. Consequently, our primary comparative analysis focuses on outdoor scenes.
Qualitative Comparison Figure 4 visually demonstrates WonderVerse’s clear superiority over existing methods. As shown in Fig 4 (a), LucidDreamer’s generated scene is far from satisfactory, blurry, obscured by an odd frame-like artifacts, and lacking in detail. WonderWorld (Figure 4 (b)), while generating a broader view, suffers from geometric discontinuities, as indicated by the visibly stitched and unnatural scene. In contrast, WonderVerse (Figure 4 (c))) produces a satisfactory and geometrically sound 3D scene. The shot of the university buildings is rendered with 3DGS [22] with high fidelity, sharp textures, coherent architecture, and natural sky and lighting. The geometric integrity is maintained seamlessly across the scene, addressing the limitations apparent in both LucidDreamer’s quality and WonderWorld’s geometric consistency. Please refer to the supplementary material for additional comparison results with existing works, including Text2room [20].
Quantitative Comparison. We quantitatively compare WonderVerse to prior work LucidDreamer [9] and WonderWorld [55], with results shown in Table 1 for both outdoor and indoor scenes. Our method demonstrates state-of-the-art scene generation, outperforming existing methods in semantic alignment, structural consistency, and perceptual quality. Notably, for outdoor scenes, WonderVerse achieves the best CLIP-S score (0.9219), indicating superior semantic alignment, along with SOTA Q-align and NIQE scores, demonstrating high generation quality. Similarly, for indoor scenes, WonderVerse significantly surpasses LucidDreamer across all metrics, achieving a new SOTA CLIP-S score of 0.9639, again affirming excellent text-scene semantic alignment and high-quality 3D indoor scene generation. Note that WonderWorld cannot generate indoor scene due to its sky layer, so we do not compare with approach in the indoor setting.
Table 1: Evaluation on novel view renderings.
Scene Type Method CLIP-S ($\uparrow$) Q-align ($\uparrow$) NIQE ($\uparrow$)
Outdoor LucidDreamer [9] 0.4200 2.6704 0.7786
WonderWorld [55] 0.7817 2.8302 0.7127
WonderVerse (ours) 0.9219 3.2229 0.7806
Indoor LucidDreamer [9] 0.4633 3.9870 0.6655
WonderVerse (ours) 0.9639 4.2680 0.7119
Figure 5: Generated 3D scene without and with our abnormal sequence detection module.
4.4 Ablation Studies
In this section, we examine the key modules of our approach, including abnormal sequence detection and scene extension. We also compare different 3D representation methods 3DGS [22] and DUSt3R [43] to demonstrate the compatibility of our approach.
4.4.1 Abnormal Sequence Detection
In this section, we study the effectiveness of our abnormal sequence detection. As illustrated in Figure 5, we present the generated 3D scene with and without abnormal sequence detection. Without our detection algorithm, noticeable artifacts degrade the generated 3D scene due to the geometric inconsistency of the generated video. In contrast, when abnormal sequences are detected and addressed, the generative quality is much higher. The improvement is further demonstrated by quantitative evaluations in Table 2. The results of both the Q-Align and NIQE metrics clearly show the significant improvements of our abnormal sequence detection module, demonstrating that this design effectively enhances the quality of the generated 3D scenes.
Table 2: Evaluation metrics of abnormal sequence detection.
Q-align NIQE
W/o abnormal sequence detection 0.67 3.39
W abnormal sequence detection 0.75 3.64
[…]
[ width=0.48autoplay=True, ]3Images/DynamicResults/NoExpand/13
[…]
[ width=0.45autoplay=True, ]3Images/DynamicResults/Expand/15
Figure 6: Video results of a generated 3D scene. Starting from (a), we extended the video to (b) with high quality and maintained good geometric consistency, resulting in a realistic extended scene (c). For the best experience of GIF animations (a, b), please view them in Acrobat or similar PDF readers.
4.4.2 Scene Extension
Figure 6 visually demonstrates how our scene extension strategy effectively increases the scale and scope of generated 3D scenes. We compare scenes generated with and without our extension technique. Figure 6 (a) shows the limited scope of the scene without extension; these images show a restricted view of the living room, focusing on a sofa section and the fireplace. In contrast, Figure 6 (b,c) vividly illustrates the expanded scene. The view is now significantly wider, revealing a larger portion of the living room, including more of the L-shaped sofa, extended window areas, and a broader view of the floor. These visuals confirm that our scene extension not only increases the scene’s scale but also successfully maintains both geometric and stylistic consistency, as evidenced by the consistent style and geometry of the furniture and room elements across the expanded view. The extended scene is richer in details and provides a more comprehensive and immersive representation of the environment compared to the limited initial scenes.
4.4.3 Efficient vs. High-quality
Our method is highly flexible and can support a wide variety of 3D representation method, such as 3DGS [22] and DUSt3R [43]. In this section, we replace 3DGS and its associated components in our pipeline with DUSt3R to demonstrate the adaptability of our approach. Experimental results reveal that DUSt3R enables efficient generation, achieving speeds over 10× faster than 3DGS. Conversely, high-quality generation can be achieved with 3DGS, as shown in Figure 7.
[…]
[ width=0.47autoplay=True, ]3Images/DynamicResults/Chris/14
Figure 7: Generated 3D scene with different 3D reconstruction methods. (a) efficient reconstruction by DUSt3R; (b) High-quality generation by 3DGS. (b) is a GIF animation, viewable in PDF readers like Acrobat.
5 Conclusion and Limitation
In this paper, we introduced WonderVerse, a framework that offers a surprisingly simple yet remarkably effective approach to generating extendable 3D scenes. Moving beyond the complexities of iterative depth estimation and image inpainting common in prior works, WonderVerse leverages the inherent world-level priors of video generative foundation models to achieve highly immersive and geometrically consistent 3D environments. Our contributions extend beyond the initial scene generation with a new technique for controllable scene expansion, enabling substantial scaling of generated environments, and an innovative abnormal sequence detection module that utilizes camera trajectory to effectively address geometric inconsistencies. As demonstrated through extensive experiments, WonderVerse can generate extendable and highly realistic 3D scenes with an elegant and streamlined pipeline, markedly outperforming existing works with more complex architectures.
Our WonderVerse framework, while effective, has certain limitations. First, The quality of our simulated 3D environments is inherently limited by existing video generation models. Second, our model mainly focuses on static 3D scene generation due to the use of 3DGS and DUSt3R for reconstruction. To overcome this, an exciting avenue for future research is combine with recent advancements in dynamic scene representation [13, 54, 46] to pave the way for generating dynamic and animated 3D environments.
Supplementary Material
6 Additional Comparisons with Existing Works
In this section, we provide additional comparison results with LucidDreamer [9], WonderWorld [55], and Text2Room [20]. As shown in the Fig 89, As visually demonstrated in Figures 8 and 9, our WonderVerse framework can generate extendable 3D scenes, achieving a significant improvement in quality, plausibility, and geometric coherence compared to existing methods.
Baseline approaches that rely on discrete image generation pipelines exhibit noticeable geometric inconsistencies. For instance, LucidDreamer produces fragmented and discontinuous 3D scenes, while Text2Room results in largely distorted 3D scenes, both exhibiting significant geometric inconsistency. WonderWorld, while improved, still exhibits discontinuities between subscenes generated iteratively, which compromises its overall visual quality and realism. On the contrary, WonderVerse produces extendable 3D scenes that are not only visually superior but also maintain geometric integrity, resulting in more believable and immersive 3D environments. This visual comparison underscores the advantage of our approach in creating extendable and coherent 3D scenes.
Figure 8: Qualitative comparison with existing works for extendable 3D scene generation.
Figure 9: Qualitative comparison with existing works for extendable 3D scene generation.
7 Additional Qualitative Results
We show additional qualitative results in Fig 11 13 15 17. We also provide the visualization of their multi-viewed point clouds as shown in Fig 10 12 14 16.
As shown in Figure 10, the reconstructed point clouds from two viewpoints capture the structural elements of an art studio, including walls, windows, and scattered objects. Figure 11 then showcases the rendered scene from multiple perspectives, effectively portraying a vibrant art studio environment populated with canvases and art supplies, consistent with the input text prompt.
Similarly, Figure 12 visualizes the 3D point cloud representing the garden’s layout, capturing the fountain and surrounding floral arrangements. Figure 13 demonstrates the rendered scene, which vividly realizes a colorful flower garden with a central fountain reflecting the sky, faithfully adhering to the descriptive text.
Besides, Figure 14 shows the point cloud reconstruction of the Taj Mahal and its surrounding environment, clearly outlining the iconic structure and pool. Figure 15 showcases the rendered scene, capturing the majestic Taj Mahal with its reflecting pool and cypress trees, demonstrating our framework’s ability to generate complex architectural scenes from textual descriptions.
Beyond generating realistic 3D scenes, our WonderVerse demonstrates the ability to create imaginative scenes that do not exist in the real world, and in diverse styles. As seen in Figure 16, the point cloud reconstruction effectively captures the intricate structure of a Lego space station, including distinct modules, astronauts, and spacecraft elements against a backdrop of stars. Figure 17 then presents the rendered scene, vividly bringing to life a futuristic Lego space station in orbit. This example, rendered in a distinct LEGO style, illustrates our framework’s capability to move beyond photorealistic scene generation. It demonstrates that WonderVerse can interpret textual descriptions to create stylized and fictional 3D environments, expanding its creative potential beyond realistic simulations.
Across these diverse examples, our method consistently generates geometrically plausible and visually compelling 3D scenes that align well with the provided text prompts.
Figure 10: Reconstructed point clouds of our WonderVerse.
Figure 11: Rendered scene of our WonderVerse.
Figure 12: Reconstructed point clouds of our WonderVerse.
Figure 13: Rendered scene of our WonderVerse.
Figure 14: Reconstructed point clouds of our WonderVerse.
Figure 15: Rendered scene of our WonderVerse.
Figure 16: Reconstructed point clouds of our WonderVerse.
Figure 17: Rendered scene of our WonderVerse.
Figure 18: WonderVerse’s interactive generation (a), guided by new text prompts, creates richer 3D scenes compared to non-interactive generation (b), which uses a single initial prompt for all subscenes.
8 Interactive Scene Generation
WonderVerse also enables interactive 3D scene generation. Starting with an initial video based on a text prompt, users can interactively guide scene extension by providing new text prompts to add desired objects. For example, as shown in Figure 18 (a), we interactively extended the scene by adding “A huge fountain” on the right side and “A huge library” on the left side, both successfully incorporated into the 3D scene. Compared to Figure 18 (b), where the entire scene is generated from a single initial prompt, the interactive approach in Figure 18 (a) produces a richer scene aligned with user’s prompts. This example showcases how WonderVerse empowers users to interactively create their own customized and expandable 3D scenes.
9 Details on Efficient 3D Reconstruction with DUSt3R
With DUSt3R [43] as our 3D reconstructor, we capture 25 frames of each video and feed them to DUSt3R and create point cloud with corresponding camera poses. Afterwards, we adopt ICP [3, 37] to register the sub-scene point cloud and finally output a point cloud containing the entire 3D scene.